The length of a rectangle is 7m more than its worth, and the area of the
rectangle is 98m^2. Find the dimensions of the rectangle.

Answer by stanbon(75887)

The length of a rectangle is 7m more than its width, and the area of the
rectangle is 98m^2. Find the dimensions of the rectangle.
------------
Area = (length)(width)
Perimeter=2(length + width)
-----------------------------
Let the width be x meters ; length = x+7 meters
-------------
98 = x(x+7)
x^2+7x-98 = 0
(x+14)(x-7)=0
x=7 meters (width of the rectangle)
x+7=14 meters (length of the rectangle)
===================
Cheers,
Stan H.